OpenStudy (anonymous):

how do you graph y = |x + 3| + 2

OpenStudy (anonymous):

plot a few values use x=-5 use x=-3 use x=-0 use x=3 use x=5 to see the relationship

OpenStudy (anonymous):

The above graph is incorrect! The graph of y = abs(x+3) is a V-shaped graph shifted 3 units to the left; so raise this V-shaped graph 2 units upwards..as y = abs(x-3) + 2 means to take y = abs(x-3) raised 2 units.

OpenStudy (anonymous):

when x=-5 y=? when x=-3 y=? when x=-0 y=? when x=3 y=? when x=5 y=?

OpenStudy (anonymous):

(-5, 4) (-3, 2) (0, 5) (3, 8) (5, 10)

OpenStudy (anonymous):